Find the extreme values of


x^4+y^4-2(x-y)^2

1
Expert's answer
2022-02-03T12:33:04-0500

The matrix "H(0,0)" is negative semidefinite sine

"\lang
\begin{bmatrix}
 -4 & 4 \\
 4 & -4
\end{bmatrix}
\begin{bmatrix}
 x \\
 y 
\end{bmatrix},
\begin{bmatrix}
 x \\
 y 
\end{bmatrix} 
 \rang =-4(x-y)^2\le0"

with equality if x=y

This is a necessary condition for a local maximum, but not sufficient. Therefore, the test is inconclusive.

Indeed, (0,0) is a saddle point. If we approach (0,0) along x=y then we have

"f(x,x)=2x^4>0"

However, if we approach along "x=-y" , then

"f(x,-x)=2x^4-8x^2<0"

near (0,0).

extrema at "x=(-\sqrt2,\sqrt2)" and "y=(\sqrt2,-\sqrt2)"